Evaluate the given function for the value specified.
F(x)= X(squared)-7X+10;find (a) f(0), (b)f(5)and (c)f(-2)
Rewrite the given equation as a function of x:
-3x+4y=11
Graph the given function:
f(x)= -2x-5
if f(x)=4x-3 find f(-1)

Answer by cheesey(22)

i. F(x) = x^2 - 7x + 10 substitute each value of x given.
a)f(0) = 0^2 - 7(0) + 10 = 10
b) f(5) = 5^2 - 7(5) +10 = 25 - 35 + 10 = 0
c) f(-2) = (-2)^2 - 7(-2) + 10 = 4 + 14 + 10 = 28
ii. Rewrite -3x + 4y = 11 as a function of x. You must solve the given equation for y, since y = f(x).
then (3x + 11)/4 = y gives f(x) = (3x + 11)/4.
iii. Graph the function using the intercepts : y-intercept (x= 0) is y = -2(0) -5 = -5. So one point on the line is (0, -5). The x-intercept (y=0) is 0 = -2x - 5 => 2x = 5 => x = 5/2.
iv. If f(x) = 4x - 3, find f(-1). Use the same technique as (i) above. Thus f(-1) = 4(-1) - 3 = -4-3 = -7.